OpenStudy (anonymous):

[4_07]Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4.

OpenStudy (cwrw238):

the required equation will have the same slope as y = -3x + 4 whose slope is -3 (the number multiplying x) so using the general form y - y1 = m(x - x1) where m = slope and (x1,y1) is any point on the line we get: y - 6 = -3(x - 2) simplify this to get your required equation

OpenStudy (anonymous):

Thank you I have answered this wrong twice already. Can you help me simplify this? step by step?

OpenStudy (cwrw238):

* y = -3x - 4

OpenStudy (cwrw238):

y - 6 = -3(x - 2) y - 6 = -3x + 6 y = -3x + 6 + 6 y = -3x + 12

OpenStudy (anonymous):

Thats the answer I got and its wrong. The choices are y=-3x, y=-3x+12 and y=3x. thank you for helping

OpenStudy (cwrw238):

thats the right answer

OpenStudy (anonymous):

thank you, im writing down the steps. Yeah!

OpenStudy (anonymous):

You are right! I added a negative.